find the center, faci, and vertices of the ellipse and sketch its graph.
x2+2x+2y2+8+1=0

Answer by lwsshak3(11628)

find the center, faci, and vertices of the ellipse and sketch its graph.
x2+2x+2y2+8+1=0
complete the squares
(x^2+2x+1)+2y2=-9+1
(x+1)^2+2y^2=-8
no solution, this is not a valid equation